OpenStudy (anonymous):

Differentiate y= (2x-5)^4 (8x^2-5)^-3 I've gotten as far as -3(2x-5)^4 (8x^2-5)^-4 (16x) + 4(8x^2-5)^-3(2x-5) * 2 The answer is 8(2x-5)^3(8x^2-5)^-4(-4x^2+30x-5) How do I get the answer?

OpenStudy (amistre64):

product rule, base rule, chain rule

OpenStudy (amistre64):

totally saw that wrong

OpenStudy (amistre64):

power rules not exponent rules

OpenStudy (amistre64):

\[y = f^p~g^q\] \[y' = pf^{p-1}~f'~g^q+f^p~q~g^{q-1}~g'\]

OpenStudy (amistre64):

\[ y= (2x-5)^4~~(8x^2-5)^{-3}\] \[ y= 4(2x-5)^3~f'~~(8x^2-5)^{-3}-3(2x-5)^4~~(8x^2-5)^{-4}~g'\] \[ y= 4(2x-5)^3~(2)~~(8x^2-5)^{-3}-3(2x-5)^4~~(8x^2-5)^{-4}~(16x)\]

OpenStudy (amistre64):

to clean it up, that looks to arrange as:\[y= 8\left(\frac{2x-5}{8x^2-5}\right)^3~-~48x~\left(\frac{2x-5}{8x^2-5}\right)^4\]

OpenStudy (amistre64):

you simply have to keep better track of whats going on; dont try to formulate it all at once, but work it out in pieces

OpenStudy (anonymous):

so your answer is somehow the same as my book answer